OpenStudy (anonymous):

What is the quotient of five over eighteen divided by five over twelve in simplest form? one over three two over three seventeen over twenty-three twenty-five over two hundred sixteen HELPPP

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

WAT DO I DO THOMASTER

thomaster (thomaster):

\(\large\sf\dfrac{~\dfrac{5}{18}~}{\dfrac{5}{12}}\)

OpenStudy (anonymous):

SO WHAT DO I DO

OpenStudy (texaschic101):

5/18 divided by 5/12 do you know how to divide fractions ?

OpenStudy (anonymous):

no

thomaster (thomaster):

\(\large\sf\dfrac{~\dfrac{5}{18}~}{\dfrac{5}{12}}\to\dfrac{5}{18}*\dfrac{12}{5}\to\dfrac{5*12}{18*5}\to \dfrac{60}{90}\)

OpenStudy (texaschic101):

flip what you are dividing by, then multiply.....

thomaster (thomaster):

you turn the second fraction upside down

thomaster (thomaster):

then multiply and solve. You can simplify 60/90 to 2/3

OpenStudy (texaschic101):

now remember to reduce the fraction

OpenStudy (anonymous):

watz the answer

OpenStudy (texaschic101):

2/3

OpenStudy (anonymous):

fan me guys